OpenStudy (nannie):

Show me how to solve this, PLEASE! 8/v=9/v-1/3

OpenStudy (anonymous):

is this \[\frac{8}{v}=\frac{9}{v}-\frac{1}{3}\]?

OpenStudy (anonymous):

is so, you can rewrite as \[\frac{1}{3}=\frac{9}{v}-\frac{8}{v}\] and since the denominators are the same, subtracting is easy. you get \[\frac{1}{3}=\frac{1}{v}\] now take the reciprocal of both sides (flip them) to get \[3=v\]